OpenStudy (anonymous):

The surface area of a sphere is 200 cm squared. If the radius were three times as large, what would the surface area be? Can you explain step by step?

OpenStudy (anonymous):

just do width times hieght by 3

OpenStudy (anonymous):

wait never mind thats not right

OpenStudy (anonymous):

ummm ohhhhh chet idunno sorry steph

OpenStudy (anonymous):

surface aread of a sphere is given by \[4*\pi*r^2\] you already have the area given.

OpenStudy (anonymous):

but at least yuh got a 1000 pple to help:)

OpenStudy (anonymous):

so \[200 = 4*\pi*r^2\]now what will happen if we had a radius 3x as large?

OpenStudy (anonymous):

\[200 = 4*\pi*r^{2}\] \[r = \sqrt{\frac{ 200 }{ 4\pi }}\] You have r, now multiply both sides by 3 \[3r = 3*\sqrt{\frac{ 200 }{ 4\pi }}\] Put it back into the formula \[(3*\sqrt{\frac{ 200 }{ 4\pi }})^{2}*4*\pi\]

OpenStudy (anonymous):

Thanks! I don't have a calculator with me, though. :/

OpenStudy (anonymous):

Nice one @DoritosHero good job :)